[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Tue Aug 29, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed Aug 30 03:13:57 UTC 2017


HSU Shift Log for Tue Aug 29 2017 23:13:50 GMT-0400 (EDT)

Shift: R

Forecaster: J Cangialosi

Technical Issues Summary:

ATCF issues:

N-AWIPS issues: Tuna froze up shortly after the day shift left. Monica  
reboot the system and it seems to be working ok now.

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: Issue with finalize hazard grid at 03z caused a delay  
in issuance of the TCV. Notified Pablo who talked Cody through procedure to  
revert the grid. Sent about 10 after.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170830/9276d3a3/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list